to form an open box, squares of side 4 cm are cut from each corner of a square piece of tin and then the sides of the tin are turned up. If the volume of the box is to be 400 cubic centimeter. What should be the area of the original piece of tin be?

Answer by cleomenius(959)

(x-4)^2 *4 = 400 We can divide each side by 4.
(x-4)^2 = 100
x^2 -8x +16 = 100. minus 100 from each side.
x^2 - 8x -84 = 0
(x -14)(x+6)
14, -6.
disregard the -6.
(10)(10)(4) = 400 cm cubed.
The original tin was 14 cm by 14 cm, area = 196 cm^2.
Cleomenius.